OBSERVATION AND ANALYSIS OF LINE WIDTH OF Q_{1}(1) IMPURITY PAIR SPECTRUM IN SOLID PARAHYDROGEN

\underline{YU ZHANG}, and TAKESHI OKA, Department of Chemistry, and Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL 60637.

The Q_{1}(1) (v = 1 ← 0, J = 1 ← 1) rovibrational spectrum of the J = 1 ortho-H_{2} impurities in solid para-H_{2}, recorded in our laboratory using a color center laser spectrometer, shows a variety of line widths, spanning nearly two orders of magnitude. The sharpest line width observed, 10 MHz HWHM for 0.2% o-H_{2} at around 4.2 K, belongs to certain Q type transitions, which preserve the F, M quantum number in the ground and excited state resulting in minimum inhomogeneous broadening. Despite the tone burst modulation used in the data recording, which distorts the line shape, true line width is derived for Lorentzian line profile. Our results relate to the previous work in microwave region very well^{a}. The broadening mechanism will be discussed including phonon relaxation and inhomogeneous broadening.

aB. W. Statt, W. N. Hardy, and R. Jochemsen Can. J. Phys. \underline{58}, 1326 (1980).